# Efficient Statistical Estimation for Sequential Adaptive Experiments with Implications for Adaptive Designs

### v1

@@ 4 Asymptotic Normality
There exist $\sigma^{2}_{0}\in\mathbb{R}^{+}$ such that $P_{Q_{0},\bar{g}_{n}}D(Q_{\infty},\bar{g}_{\infty})^{2}\overset{p}{\to}\sigma^{2}_{0}.$
@@
$\bar{\mathcal{Q}}$ is a class of functions with finite sup-norm covering number $N_{\infty}({\epsilon},\bar{\mathcal{Q}})$.
@@
Suppose $\inf_{a\in\mathcal{A},w\in\mathcal{W}}\bar{g}_{\infty}\geq\eta$ for some $\eta>0$. Assume that $v_{n}(\bar{Q}_{\infty},\bar{g}_{n},\bar{g}_{\infty})=o_{P}(1)$, where $v_{n}$ is defined by
@@
where $\rho(x)=e^{x}-x-1$ and $c_{2}=8/{\eta}^{2}$.
@@
Theorem 6 (Asymptotic Normality).
@@
5 Efficiency
In this section, we provide efficiency results of ADL-TMLE for estimating various causal estimands based on data collected from adaptive experiments. We start by defining a Hilbert space $\mathcal{H}(Q,g)=\{h\in L_{2}^{0}(P_{Q,g}):h(o)=h_{w}(w)+h_{y}(w,a,y),\int h_{w}(w)dQ_{w}(w)=0,\int h_{y}(w,a,y)dQ_{y}(y|a,w)=0\}$ for any $Q=(Q_{W},Q_{Y})\in\mathcal{Q}$, with a referencing treatment randomization distribution $g\in\mathcal{G}$ and the corresponding density $p_{Q,g}(o)=q_{w}(w)g(a|w)q_{y}(y|a,w)$ under $P_{Q,g}$. For any $h_{1},h_{2}\in\mathcal{H}(Q,g)$, define their inner product as $h_{1},h_{2}>_{\mathcal{H}(Q,g)}:=\int h_{1}(o)h_{2}(o)dP_{Q,g}(o)$. Define a class of paths $\{Q^{h}_{\epsilon}=(Q_{w,\epsilon}^{h},Q_{y,\epsilon}^{h}):dQ_{w,\epsilon}^{h}=(1+\epsilon h_{w})dQ_{w},dQ_{y,\epsilon}^{h}=(1+\epsilon h_{y})dQ_{y}\}$. We define the fluctuated distribution $P^{n}_{Q_{\epsilon_{n}}^{h}}$ as one with density $p^{n}_{Q_{\epsilon_{n}}^{h},g_{1},\cdots,g_{n}}=\prod_{i=1}^{n}q_{\epsilon_{n}}^{h}(o_{i})g_{i}(o_{i})$, where $\epsilon_{n}=1/\sqrt{n}$.
@@ 5.2 Regularity and Efficiency
Now we introduce the definition of regular estimator and the limit distribution of regular estimators of regular parameters in asymptotically normal experiments based on the convolution theorem. Define a sequence of parameters $\psi(Q_{\epsilon_{n}}^{h}):=E_{Q_{w,\epsilon_{n}}^{h}}[\bar{Q}_{y,\epsilon_{n}}^{h}(1,W)-\bar{Q}_{y,\epsilon_{n}}^{h}(0,W)]$ for every $h\in\mathcal{H}(Q,\bar{g}_{\infty})$, where $Q_{w,\epsilon_{n}}^{h}$ denotes the perturbed distribution of $Q_{W}$ and $\bar{Q}_{y,\epsilon_{n}}^{h}$ is the conditional mean function of the outcome under the perturbed distribution of $Q_{Y}$.
@@
for a continuous, linear map $\dot{\psi}:\mathcal{H}(Q,\bar{g}_{\infty})\to\mathbb{R}$ and a certain norming operator $\sqrt{n}$. Here, $\dot{\psi}(Q)(h)=<D(Q,\bar{g}_{\infty}),h>_{\mathcal{H}(Q,\bar{g}_{\infty})}$ is the pathwise derivative of $\psi(Q_{\epsilon_{n}}^{h})$ at $\epsilon=0$.
@@
Theorem 8 (Limit Distribution of Regular Estimators by Convolution Theorem).
@@
Suppose Assumptions 3, 5, 6, 7, 8, 9 and 10 hold with $Q_{\infty}=Q_{0}$, then $\psi(\bar{Q}_{n}^{*})$ is efficient in that the asymptotic variance of $\sqrt{n}(\psi(Q_{n}^{*})-\psi(Q_{0}))$ achieves the minimal possible variance $\|\dot{\psi}^{*}(Q_{0})\|_{\mathcal{H}(Q_{0},\bar{g}_{\infty})}^{2}$ of the limit distribution of regular estimators of $\psi(Q_{0})$ under the asymptotically normal experiments $(\mathcal{\bar{O}}(n),\mathcal{F}_{n},P^{n}_{Q_{\epsilon_{n}}^{h}}:h\in\mathcal{H}(Q_{0},\bar{g}_{\infty}))$.
@@
Second, for AD-TMLE $\psi(\bar{Q}_{n}^{\dagger})$ to be asymptotically normal, $g_{n}$ needs to stabilize such that $\frac{1}{n}\sum_{i=1}^{n}P_{Q_{0},g_{i}}D(Q_{\infty},g_{i})^{2}$ stabilizes as $n\to\infty$. In contrast, our proposed ADL-TMLE $\psi(\bar{Q}_{n}^{*})$ requires only the average design $\bar{g}_{n}$ to stabilize. Likewise, while the efficiency proof for ADL-TMLE can be extended to AD-TMLE, the latter requires stronger positivity and convergence conditions on the adaptive design than those needed for ADL-TMLE.
@@ 6 Adaptive Designs Motivated by Efficiency Results
For a given causal estimand of interest, one can characterize an oracle experimental design that minimizes semiparametric efficiency bound for estimating that estimand using the data collected through that oracle design, which we denote by $\pi^{*}:\mathcal{W}\times\mathcal{A}\to(0,1)$. Using ATE as an example, it is known that the ideal oracle design $\pi^{*}$ that minimizes the semiparametric efficiency bound follows the conditional Neyman allocation: $\pi^{*}(a|w):=\frac{\sigma(a,w)}{\sigma(1,w)+\sigma(0,w)},$ where $\sigma^{2}(a,w):=Var[Y|A=a,W=w]$.
However, the oracle design is unknown in practice. This necessitates adaptive experimental designs that sequentially learn the oracle design from the accumulated data and determine treatment randomization for newly enrolled units according to its evolving estimate, with the aim of improving estimation efficiency. Motivated by efficiency results of ADL-TMLE for adaptive experiment data and its connection to a reference i.i.d. experiment characterized by the average design, we propose in this section a novel adaptive design that guides the average of adaptive design toward the oracle design, which we refer to $\bar{g}$-driven adaptive design”.
The $\bar{g}$-driven adaptive design is constructed as follows. At the beginning, we enroll $n_{0}$ units in the experiment and randomly assigning treatment with a baseline probability (e.g. 0.5). Here, $n_{0}$ is a user-specified “burn-in” sample size to estimate Neyman allocation for adapting treatment randomization probabilities. After that, suppose we have enrolled $i$ units. Denote $\bar{g}_{i}=\frac{1}{i}\sum_{j=1}^{i}g_{j}$ as the current average of adaptive design. Upon the enrollment of each new unit $i+1$, we use the accumulated data $\bar{O}(i)=(O_{1},\dots,O_{i})$ to construct $\hat{\pi}_{i+1}^{*}$ as an estimate of the oracle design $\pi^{*}$ for the causal estimand of interest. In the example with the ATE as the target, $\hat{\pi}_{i+1}^{*}$ is constructed by first estimating the unknown conditional variance $\sigma^{2}(a,w)$ by $\hat{\sigma}^{2}_{1:i}(a,w)$, and then plugging that into the Neyman allocation formula: $\hat{\pi}_{i+1}^{*}(a|w)=\frac{\hat{\sigma}_{1:i}(a,w)}{\hat{\sigma}_{1:i}(1,w)+\hat{\sigma}_{1:i}(0,w)}.$ We then integrate this estimate into $\overline{\hat{\pi}}_{i+1}^{*}:=\frac{1}{i-n_{0}}\sum_{j=n_{0}+1}^{i+1}\hat{\pi}^{*}_{j}$ as the target for the next-step average design, which mitigates the variability inherent in adapting solely to the latest estimate and produces a smoother, more stable design trajectory in finite samples. Finally, by solving $\bar{g}_{i+1}=\overline{\hat{\pi}}^{*}_{i+1}$ for the next average design, the treatment randomization function $g_{i+1}$ is given by
@@
Remark 2 (Comparison with Standard Neyman Allocation).
Compared with the adaptive designs with standard Neyman allocation that directly sets $g_{i+1}=\hat{\pi}_{i+1}^{*}$, our $\bar{g}$-driven adaptive design adopts a “self-calibrating” approach: if the average design in earlier stages presents higher probability to assign treatment (or control) relative to the estimated optimal design, the $\bar{g}$-driven adaptive design compensates in subsequent stages by reducing treatment (control) randomization probability accordingly. This design strategy offers global control over the adaptive design as a whole, steering the overall average design toward the oracle design informed by the efficiency results in a stable manner, rather than optimizing $g_{i+1}$ alone in isolation without considering previous treatment randomization in the adaptive experiment. Together, the $\bar{g}$-driven design framework and efficient estimation framework of ADL-TMLE work jointly to improve estimation efficiency in answering various causal questions of interest.
@@ 7 Simulation Study
This section presents simulation studies to evaluate the performance of the proposed estimator and adaptive design to minimize variance in ATE estimation.
@@ 7.1 Setup
Figure 1 shows the conditional expectation function and variance functions against baseline covariates in the simuation, while Figure 2 displays the optimal Neyman allocation and CATE functions. The oracle Neyman allocation function varies across covariates, resulting in heterogeneous treatment randomization probabilities in Neyman allocation. The CATE function is also heterogeneous and follows a different pattern from that of the Neyman allocation. Consequently, the benefit-driven design will result in higher variance in the ATE estimate than under an adaptive design with Neyman allocation.
@@
Our simulation study investigates two questions: (A) Estimation: Given data collected under the same type of adaptive design, we evaluate the performance of ADL-TMLE in estimating ATE and compared that with AD-TMLE. Specifically, we compare their performance under both standard Neyman-allocation-based adaptive design that aims at efficient estimation of ATE and the benefit-driven design that does not target estimation efficiency. (B) Design: With the goal of minimizing variance in ATE estimation, we evaluate the performance of the proposed $\bar{g}$-driven design and compare that with the standard Neyman allocation, oracle Neyman allocation design and the non-adaptive design.
@@ 7.2 Results
(A) Estimation. Table 1 presents the bias, variance, and MSE and the coverage rate of the 95% Wald-type confidence interval of the proposed ADL-TMLE. It also reports the coverage of the 95% Wald-type “oracle confidence interval” based on the estimator’s Monte Carlo standard deviation of the point estimates, whose nominal coverage supports the estimator’s asymptotic normality.
@@
Table 1: Performance of ADL-TMLE in adaptive experiements with benefit-driven design and standard Neyman-allocation design at time points 2, 4, 6, 8, and 10. The table includes bias, variance, mean squared error (MSE), coverage probability of 95% Wald confidence intervals based on estimated standard errors, along with coverage probability of the 95% Wald “oracle” intervals constructed by the Monte Carlo standard deviation of the point estimates.
Figure 3 compares the relative variance of the proposed TMLE estimator, $\psi(\bar{Q}_{n}^{*})$ using $\bar{g}_{n}$, with that of the TMLE, $\psi(\bar{Q}_{n}^{\dagger})$, using $g_{i}$. The results show that ADL-TMLE $\psi(\bar{Q}_{n}^{*})$ significantly reduces variance compared with AD-TMLE in adaptive experiments, under both benefit-driven design and standard Neyman allocation. This validates ADL-TMLE’s efficiency across various designs, whether aimed at the efficient estimation of causal estimand of interest or targeted for other objectives.
@@
(B) Design. Figure 4 presents the relative variance of ADL-TMLE under various designs compared with that in a non-adaptive design with a fixed treatment probability of 0.5. The results show that both adaptive designs with standard Neyman allocation and $\bar{g}$-driven Neyman allocation yield lower variance than the non-adaptive design. Moreover, the proposed $\bar{g}$-driven design achieves greater variance reduction than the standard Neyman allocation and approaches closer to the oracle design.
@@
Additionally, we evaluate the oracle versions of different adaptive designs with Neyman allocations to compare their effects on the semiparametric efficiency. Specifically, for $i\geq n_{0}$, these adaptive designs construct Neyman allocations using the true conditional variance under $Q_{0,Y}$ in place of their estimates when updating treatment randomization probabilities. Figure 5 presents the trajectory of the relative value of $D(Q_{0},\bar{g}_{n})^{2}$ over time in each type of adaptive design, indicating how quickly the asymptotic variance of ADL-TMLE approaches the semiparametric efficiency bound minimized by the oracle Neyman allocation. The results show that this quantity under the $\bar{g}$-driven design converges to that bound more rapidly than the standard Neyman allocation, supporting the advantage of the proposed $\bar{g}$-driven adaptive design of guiding the global average design toward the optimal target.

### v2

@@ 4 Asymptotic Normality
There exists $\sigma^{2}_{0}\in\mathbb{R}^{+}$ such that $P_{Q_{0},\bar{g}_{n}}D(Q_{\infty},\bar{g}_{\infty})^{2}\overset{p}{\to}\sigma^{2}_{0}.$
@@
$\bar{\mathcal{Q}}$ is a class of functions with finite sup-norm covering number $N_{\infty}({\epsilon},\bar{\mathcal{Q}})$, with $\mathcal{J}_{\infty}\left({\epsilon},{\bar{\mathcal{Q}}}\right):=\int_{0}^{\epsilon}\sqrt{\log\left(1+\mathcal{N}_{\infty}\left({u},{\bar{\mathcal{Q}}}\right)\right)}du$ the $\epsilon$-entropy integral in supremum norm of $\bar{\mathcal{Q}}$.
@@
Suppose $\inf_{a\in\mathcal{A},w\in\mathcal{W}}\bar{g}_{\infty}(a|w)\geq\eta$ for some $\eta>0$. Define
@@
where $\rho(x)=e^{x}-x-1$ and $c_{2}=8/{\eta}^{2}$. Assume that $v_{n}(\bar{Q}_{\infty},\bar{g}_{n},\bar{g}_{\infty})=o_{P}(1)$.
@@
Theorem 6 (Asymptotic Normality of ADL-TMLE).
@@
5 Semiparametric Efficiency
In this section, we provide semiparametric efficiency results of ADL-TMLE for adaptive experiments. We start by defining a Hilbert space: $\mathcal{H}(Q,g):=\{h\in L_{2}^{0}(P_{Q,g}):h(o)=h_{w}(w)+h_{y}(w,a,y),\int h_{w}(w)dQ_{w}(w)=0,\int h_{y}(w,a,y)dQ_{y}(y|a,w)=0\}$ for any $Q=(Q_{W},Q_{Y})\in\mathcal{Q}$, with a reference treatment randomization distribution $g\in\mathcal{G}$ and the corresponding density $p_{Q,g}(o)=q_{w}(w)g(a|w)q_{y}(y|a,w)$ under $P_{Q,g}$. For any $h_{1},h_{2}\in\mathcal{H}(Q,g)$, define their inner product as $h_{1},h_{2}>_{\mathcal{H}(Q,g)}:=\int h_{1}(o)h_{2}(o)dP_{Q,g}(o)$. Define a class of paths:
$\{Q^{h}_{\epsilon}=(Q_{w,\epsilon}^{h},Q_{y,\epsilon}^{h}):dQ_{w,\epsilon}^{h}=(1+\epsilon h_{w})dQ_{w},dQ_{y,\epsilon}^{h}=(1+\epsilon h_{y})dQ_{y},h\in\mathcal{H}(Q,\bar{g}_{\infty})\}$
for any $Q\in\mathcal{Q}$. We denote by $P^{n}_{Q_{\epsilon_{n}}^{h}}$ the fluctuated distribution of $P^{n}_{Q}$ with density $p^{n}_{Q_{\epsilon_{n}}^{h},g_{1},\cdots,g_{n}}$, where $\epsilon_{n}=1/\sqrt{n}$.
@@ 5.2 Regularity and Efficiency
Now we introduce the definition of regular estimator and present the limit distribution for regular estimators of regular parameters based on the convolution theorem. Define a sequence of parameters $\psi(Q_{\epsilon_{n}}^{h}):=E_{Q_{w,\epsilon_{n}}^{h}}[\bar{Q}_{y,\epsilon_{n}}^{h}(1,W)-\bar{Q}_{y,\epsilon_{n}}^{h}(0,W)]$ for every $h\in\mathcal{H}(Q,\bar{g}_{\infty})$, where $Q_{w,\epsilon_{n}}^{h}$ denotes the perturbed distribution of $Q_{W}$ and $\bar{Q}_{y,\epsilon_{n}}^{h}$ is the conditional mean function of the outcome under the perturbed distribution of $Q_{Y}$.
@@
for a continuous, linear map $\dot{\psi}:\mathcal{H}(Q,\bar{g}_{\infty})\to\mathbb{R}$ and a certain norming operator $\sqrt{n}$. Here, $\dot{\psi}(Q)(h)$ is the pathwise derivative of $\psi(Q_{\epsilon}^{h})$ at $\epsilon=0$ and equals $D(Q,\bar{g}_{\infty}),h>_{\mathcal{H}(Q,\bar{g}_{\infty})}$ by Riesz representation.
@@
Theorem 8 (Limit Distribution of Regular Estimators via the Convolution Theorem).
@@
Suppose Assumptions 3, 5, 6, 7, 8, 9 and 10 hold with $Q_{\infty}=Q_{0}$, then $\psi(\bar{Q}_{n}^{*})$ is regular and efficient, i.e., the asymptotic variance of $\sqrt{n}(\psi(Q_{n}^{*})-\psi(Q_{0}))$ achieves the minimal asymptotic variance $\|\dot{\psi}^{*}(Q_{0})\|_{\mathcal{H}(Q_{0},\bar{g}_{\infty})}^{2}$ among all regular estimators.
@@
Second, for AD-TMLE $\psi(\bar{Q}_{n}^{\dagger})$ to be asymptotically normal, $g_{n}$ needs to stabilize such that $\frac{1}{n}\sum_{i=1}^{n}P_{Q_{0},g_{i}}D(Q_{\infty},g_{i})^{2}$ converges as $n\to\infty$. In contrast, our proposed ADL-TMLE $\psi(\bar{Q}_{n}^{*})$ requires only the average design $\bar{g}_{n}$ to stabilize. Likewise, while the efficiency proof for ADL-TMLE can be extended to AD-TMLE, the latter requires stronger positivity and convergence conditions on the adaptive design than those needed for ADL-TMLE.
@@ 6 Adaptive Designs Motivated by Efficiency Results
Beyond efficient estimation after the completion of a given adaptive experiment, we investigate how to design real-time adaptive experiments to further reduce the variance of the estimator for a specific causal estimand of interest. Ideally, one can characterize an oracle experimental design that minimizes semiparametric efficiency bound for estimating that estimand using the data collected through that oracle design, which we denote by $\pi^{*}:\mathcal{W}\times\mathcal{A}\to[0,1]$. Using ATE as an example, it is known that the oracle design $\pi^{*}$ follows the conditional Neyman allocation: $\pi^{*}(a|w):=\frac{\sigma(a,w)}{\sigma(1,w)+\sigma(0,w)},$ where $\sigma^{2}(a,w):=Var[Y|A=a,W=w]$.
However, the oracle design is unknown a priori in practice. This necessitates adaptive experimental designs that sequentially learn the oracle design from the accumulated data and determine treatment randomization probabilities for newly enrolled units according to the evolving estimate of the oracle design, with the aim of improving estimation efficiency. Motivated by efficiency results of ADL-TMLE, in this section we propose a novel adaptive design that guides the average of adaptive design toward the oracle design, which we refer to $\bar{g}$-driven adaptive design”.
This design is constructed as follows. At the beginning of the experiment, we enroll $n_{0}$ units and randomly assign treatments with a baseline probability (e.g. 0.5). Here, $n_{0}$ is a user-specified “burn-in” sample size to estimate the oracle design for adapting treatment randomization probabilities. After that, suppose we have enrolled $i$ units and define $\bar{g}_{i}=\frac{1}{i}\sum_{j=1}^{i}g_{j}$ as the current average of adaptive design. Upon the enrollment of each new unit $i+1$, we use the accumulated data $\bar{O}(i)=(O_{1},\dots,O_{i})$ to construct $\hat{\pi}_{i+1}^{*}$ as an estimate of the oracle design $\pi^{*}$ for the causal estimand of interest. In the example of ATE, $\hat{\pi}_{i+1}^{*}$ is constructed by first estimating the unknown conditional variance $\sigma^{2}(a,w)$ by $\hat{\sigma}^{2}_{1:i}(a,w)$ based on $\bar{O}(i)$, and then plugging that into the Neyman allocation formula: $\hat{\pi}_{i+1}^{*}(a|w)=\frac{\hat{\sigma}_{1:i}(a,w)}{\hat{\sigma}_{1:i}(1,w)+\hat{\sigma}_{1:i}(0,w)}.$ We then integrate this estimate into $\overline{\hat{\pi}}_{i+1}^{*}:=\frac{1}{i-n_{0}}\sum_{j=n_{0}+1}^{i+1}\hat{\pi}^{*}_{j}$ as the target for the next-step average design, which mitigates the variability inherent in adapting solely to the latest estimate and produces a smoother, more stable design trajectory in finite samples. Finally, by solving $\bar{g}_{i+1}=\overline{\hat{\pi}}^{*}_{i+1}$ for the next average design, the treatment randomization function $g_{i+1}$ is given by
@@
Remark 2 (Comparison with standard Neyman-allocation adaptation).
Compared with the adaptive designs with standard Neyman allocation that directly sets $g_{i+1}=\hat{\pi}_{i+1}^{*}$, our $\bar{g}$-driven adaptive design adopts a “self-calibrating” approach: if the average design in earlier stages presents higher probability to assign treatment (control) relative to the estimated optimal design, the $\bar{g}$-driven adaptive design compensates in subsequent stages by reducing the randomization probability for treatment (control) accordingly. This design strategy offers global control over the adaptive design as a whole, steering the overall average design toward the oracle design in a stable manner, rather than optimizing $g_{i+1}$ alone in isolation without considering previous treatment randomization in the adaptive experiment. Together, the $\bar{g}$-driven design framework and efficient estimation framework of ADL-TMLE work jointly to improve estimation efficiency in answering various causal questions of interest.
@@ 7.1 Setup
This section presents simulation studies to evaluate the performance of the proposed estimator and adaptive design to minimize variance in estimating ATE as a representative example. Specifically, we examine the following aspects: (A) Estimation: We evaluate the performance of ADL-TMLE in their efficiency in estimating ATE and compared that with AD-TMLE across adaptive designs that either aim to improve estimation efficiency for the causal estimand or target a different objective. (B) Design: With the goal of minimizing variance in ATE estimation, we evaluate the performance of the proposed $\bar{g}$-driven design and compare that with other alternative designs.
@@
Figure 1 displays the oracle Neyman allocation and CATE functions across baseline covariates. The oracle Neyman allocation function varies across covariates, resulting in heterogeneous treatment randomization probabilities, while CATE function is also heterogeneous and follows a different pattern. Consequently, the benefit-driven design is expected to yield higher variance in ATE estimation than the adaptive designs using Neyman allocation.
@@ 7.2 Results
(A) Estimation. Table 1 presents the bias, variance, and MSE and the coverage rate of the 95% Wald-type confidence interval of the proposed ADL-TMLE in adaptive experiements with benefit-driven design and standard Neyman-allocation design, respecitvely. It also reports the coverage of the 95% Wald-type “oracle confidence interval” based on the estimator’s Monte Carlo standard deviation of the point estimates, with its nominal coverage supporting the estimator’s asymptotic normality.
@@
Table 1: Performance of ADL-TMLE in adaptive experiements with benefit-driven design and standard Neyman-allocation design at time points 2, 4, 6, 8, and 10. The table includes bias, variance, mean squared error (MSE), coverage probability of 95% Wald-type confidence intervals based on estimated standard errors, along with coverage probability of the 95% Wald-type “oracle confidence intervals” constructed by the Monte Carlo standard deviation of the point estimates.
Figure 2 compares the relative variance of the proposed TMLE estimator ADL-TMLE $\psi(\bar{Q}_{n}^{*})$ with that of AD-TMLE $\psi(\bar{Q}_{n}^{\dagger})$. The results show that ADL-TMLE $\psi(\bar{Q}_{n}^{*})$ significantly reduces variance compared with AD-TMLE in adaptive experiments, under both benefit-driven design and standard Neyman allocation, demonstrating ADL-TMLE’s efficiency across different adaptive experiments with varying design objectives.
@@
(B) Design. Figure 3 presents the relative variance of ADL-TMLE under various designs compared with that in the non-adaptive design with fixed treatment randomization probability 0.5. The results show that both adaptive designs with standard Neyman allocation and the proposed $\bar{g}$-driven Neyman allocation yield lower variance than the non-adaptive design. Moreover, the $\bar{g}$-driven adaptive design consistently results in a smaller variance than the adaptive design with standard Neyman allocation and approaches closer to the oracle design.
@@
Additionally, we evaluate the effects of different designs on semiparametric efficiency for ATE estimation. Specifically, we construct oracle versions of Neyman-allocation-based adaptive designs, in which the treatment randomization probabilities are based on the oracle Neyman allocation functions using the true conditional variances rather than the estimated ones for units $i\geq n_{0}$. Figure 4 plots the trajectory of the relative value of $D(Q_{0},\bar{g}_{n})^{2}$ over time for these designs (including the non-adaptive one) against the ideal oracle design that applies the oracle Neyman allocation from the outset. This illustrates how quickly the variance of ADL-TMLE under an adaptive design approaches the semiparametric efficiency bound minimized by the oracle Neyman allocation. The results show that this quantity under the $\bar{g}$-driven adaptive design converges to the oracle one more rapidly than the standard Neyman allocation, highlighting the advantage of guiding the global average design toward the optimal target.